# RelCon: Relative Contrastive Learning for a Motion Foundation Model for Wearable Data

### v4

@@ 4.1 Foundation Model Pretraining
We trained models on Inertial Movement Unit (IMU) sensors from the Apple Heart & Movement Study (AHMS) (MacRae, 2021). AHMS is an ongoing research study designed to explore the links between physical activity and cardiovascular health, which is sponsored by Apple and conducted in partnership with American Heart Association and Brigham and Women’s Hospital. To be eligible for the study, participants must — among other eligibility criteria – be at least 18 years of age (at least 19 in Alabama and Nebraska; at least 21 in Puerto Rico), reside in the United States, have access to an Apple Watch, and provide informed consent electronically in the Apple Research app.
The training data included a subset of study data with 87,376 participants recorded over one day, with a 10/3/3 train/val/test split. Following prior convention (Reyes-Ortiz et al., 2015), we use 2.56 seconds of the raw 100 Hz 3-axis x,y,z accelerometry signal of the IMU sensor in the wearable device as input to our embedding model. Each model was pre-trained with 8 x A100 GPUs for 24 hours. Models iterated over a total of 1 billion samples, and a total of $\sim$30k unique participant-days.
We evaluated RelCon along with other commonly used methods in SSL with Accel for comparison:
  • Accel SimCLR (Tang et al., 2020; Chen et al., 2020) contrasts positive pairs formed by accel-specific augmentations (i.e. 3D rotation) and has been shown to have state-of-the-art performance in the accel domain (Haresamudram et al., 2022).
  • Augmentation Prediction (Yuan et al., 2024; Haresamudram et al., 2022) predicts whether an accel-specific augmentation was applied to each sample.
  • Accel REBAR (Xu et al., 2024) uses a contrastive loss where positive pairs are identified by a learned motif-similarity. Instead of using the original version of REBAR designed for generalized time-series data, we utilize it with the accel-focused innovations proposed in Section 3.1.
For each of these methods, we used a 1D ResNet-34 encoder backbone that used average pooling to generate a 256-dimensional embedding vector (3.9M parameters).
@@ 4.2 Downstream Evaluation
First, in order to assess generalizable of our learned embedding, we evaluate our models with our Task Diversity Evaluation with two different datasets, in which we compare the RelCon foundation model against a set of diverse downstream tasks and to self-supervised models trained on our data with three other methods that have seen strong performance accel data: SimCLR, AugPred, and REBAR. We use the embeddings with linear regression for various gait metrics, and with linear probe classification on both the subsequence and workout-level.
Next, in order to compare against prior work, we evaluate our RelCon foundation model with our Benchmarking Evaluation with four additional classification datasets. Specifically, we compare against another large-scale pre-trained accel foundation model (Yuan et al., 2024) and against an accel SSL benchmarking study (Haresamudram et al., 2022), with comparisons to seven distinct self-supervised learning methods. These datasets help evaluate model generalizability under data distribution shifts, including differing sensor positions and inference window lengths.
@@
In total, we compare against 6 different downstream datasets across 4 different types of downstream tasks. We compare our RelCon foundation model against 11 models total: 3 pre-trained from scratch in the Task Diversity Evaluation and 8 from the prior literature in the Benchmarking Evaluation.
4.2.1 Task Diversity Evaluation Datasets and Set-up
@@
Gait Metric Regression: We used a dataset including gait mat collection to evaluate models on gait metric regression (Apple, 2021). All participants completed proctored overground walking tasks with two mobile devices with IMU sensors at each side of the body in different locations (i.e. at the hip, in a front or back pocket, or in a waist bag). We used the Design Cohort A participants, which included 359 participants with an average age of 74.7, who provided informed consent. Participants were instructed to conduct 4 different walking tasks: 1) one lap at self-selected speed, 2) four laps at an instructed slow speed, and 3) as many laps as possible within 6 minutes. Each walking task was conducted along a 12-meter straight-line course, with an 8-meter pressure mat placed centrally and various statistics about each participant’s gait were calculated: step count, walking speed, step length, double support time, and walking asymmetry. Models were evaluated on predicting double support time (DST) and stride velocity.
@@
Each 2.56s subsequence was matched to the lap aggregated target (e.g. total number of steps or average walking speed). The participants were assigned into 50% train and 50% test randomly based on participant ID, where every lap for a given participant falls into the same split. The training split was used to train linear regression probes on embeddings from the self-supervised models. Metrics were selected following a prior report (Apple, 2021): mean squared error (MSE), std dev of squared error (SDSE), mean absolute error (MAE), std dev of absolute error (SDAE), and Pearson’s Correlation Coefficient. Mean and std devs were calculated by aggregating predictions across all inputs of a given user and are related to bias and variance respectively. Correlation is used in order to assess how each user’s average gait metric corresponds to ground truth values. Ranges for each metric were calculated by retraining the linear regression probe five times with different set seeds.
@@
Activity Classification: We evaluated activity classification performance using self-reported activity labels gathered from data in AHMS. We used a subset of data with $\sim$2k total users across 14 workouts (full list in Table 2). The 14 selected workouts captured a range of diverse activities (e.g. kickboxing and rowing) that are non-trivial to separate (i.e. outdoor cycling vs. indoor cycling). For evaluation, workouts were class balanced so each included $\sim$22 hours of data, for a total of 310 workout hours. We used a 4/1/5 train/val/test split based on participant ID. We ensured the subset of data used for classification did not overlap with the pre-training dataset, preventing any data leakage.
@@
Embeddings are generated for each 2.56s-long subsequence. We evaluate classification on both the subsequence-level as well as the workout-level. At the workout-level, we predict workout classes across each of 2.56s subsequences within the workout, and select the most frequently predicted workout. Including two prediction scales allow us to evaluate provides additional context about how information learned by the embedding interacts with time aggregation. We report F1, Kappa, Accuracy, and macro AUC metrics following previous work (Haresamudram et al., 2022; Yuan et al., 2024; Xu et al., 2024). Ranges per metric are obtained by retraining the probe five times and calculating the mean and standard deviation.
4.2.2 Benchmarking Evaluation Datasets and Set-up
@@
Comparison to a Large-Scale Pre-trained Accel Model: Yuan et al. (2024) proposes an Accel foundation model with large-scale pre-training on the UK BioBank (Bycroft et al., 2018) dataset. They are able to train over $\sim$100k unique participants for a total of $\sim$700k unique participant-days. Their foundation model is based upon an augmentation prediction approach with a ResNet-18 backbone, and they use it to train and embed 10s subsequences of a 3-axis accelerometry signal collected at the wrist. Please refer to the original work for more details (Yuan et al., 2024).
@@
Following their evaluation methodology, we directly compare our RelCon foundation model against their model along three different dimensions: fine-tuning the entire model, using an MLP probe, and training from scratch. We exactly mimic their cross-validation splits on the Opportunity (Roggen et al., 2010) and PAMAP2 (Reiss, 2012) activity classification datasets and use these splits to generate the metric ranges. We also match their 10s length used for inference.
@@
Comparison to an Accel SSL Benchmarking Study: Haresamudram et al. (2022) seeks to assess the current state of the accelerometry self-supervised learning field, in the context of activity classification. To this end, they benchmark a diverse range of self-supervised, including our aforementioned accel-specific methods, Accel SimCLR (Tang et al., 2020) and Augmentation Prediction (Saeed et al., 2019), as well as generalized SSL methods, such as SimSiam (Chen & He, 2021) and BYOL (Grill et al., 2020). Each of their SSL approaches have a unique backbone architecture that corresponds to their original work. They pre-train their SSL methods on the 3-axis wrist-mounted accelerometry signals from Capture-24 dataset (Chan Chang et al., 2021), which is composed of 151 unique participants for a total of $\sim$4k participant-hours.
@@
Similar to our RelCon method, each of their SSL methods are trained on wrist data, but for evaluation, they use a different activity classification datasets that each correspond to accelerometry signals collected at different positions. We compare our RelCon FM with a downstream dataset at each benchmarked position: HHAR (Blunck et al., 2015) for Wrist, Motionsense for Waist (Malekzadeh et al., 2018), and PAMAP2 (Reiss, 2012) for leg. In this way, each of their methods, as well as our RelCon FM, will have pre-trained on wrist accelerometry data and then evaluate on potentially different sensor position. We adopt the same exact cross-validation splits in the original work and use them to generate our metric ranges. We match their 2s length used for inference.
@@ 5.1 Task Diversity Evaluation Results
Table 1: Results of Gait Metric Regression. Mean and standard deviations of Squared Error and Absolute Error were calculated by aggregating predictions across given user and are related to bias and variance respectively. Strong consistent performance of RelCon shows that our model is able to do well on capturing user-specific information for regression.
Figure 3: Plot of correlation between predicted and true DST. RelCon has the highest correlation.
Gait metric regression: Table 1 shows gait metric regression performance of each SSL method. RelCon had the strongest consistent performance across both gait metrics, double support time and stride velocity. While methods based on prior work with Accel SimCLR (Haresamudram et al., 2022; Tang et al., 2020) were not developed with a focus on gait, they still achieve strong performance, particularly for stride velocity. Although REBAR also uses a learnable distance, it generally has lower performance than RelCon. RelCon’s relative constrastive loss seems to improve the model’s ability to capture fine-scale differences between subsequences of differing walking speeds or DSTs.
AHMS Classification (Wrist→Wrist)
@@
Table 2: Results on Activity Classification. RelCon achieves consistently strong performance when evaluated both at the subsequence-level and workout-level. SimCLR does well at the subsequence-level, but performs much worse at the workout-level. The opposite is true for AugPred. This show that there exist nuances captured by our models at the subseq- vs. workout-level.
Activity classification: Table 2 summarizes activity classification results on the activity classification dataset described in Sec. 4.2.1. RelCon embeddings had distinct cluster separability, highlighted in the its t-SNE visualization in Fig. 2, and further detailed confusion matrics can be found in Appendix A.2. SimCLR particularly had low performance in workout-level prediction of stair climbing, which was frequently confused as climbing (Fig. 4 and 5). SimCLR has worse performance in workout-level classification metrics compared to the subseq-level ones, and the opposite is true for AugPred. RelCon is consistently strong across both subseq and workout-level. RelCon is better at classifying stair climbing compared to the other methods, which will often confuse stair climbing with climbing or elliptical (Fig. 4 and 5). These three classes have slower, deliberate hand swinging motions, but we hypothesize that RelCon’s resistance to false positives and negatives enables the model to capture subtle differences. Additionally, RelCon can better predict outdoor running from indoor running. We hypothesize that RelCon can better identify when the running pattern is more uniform, implying that the running is perhaps being done indoors.
Figure 4: ROC curves of AHMS Workout Level Classification Task. Amongst other performance improvements, unlike the other approaches, RelCon is able to clearly better classify stair climbing.
@@ 5.2 Benchmarking Evaluation Results
RelCon FM MLP Probe AHMS 69.1 ± 8.3 62.1 ± 6.2 85.38 ± 3.6 84.73 ± 3.6
@@
RelCon FM Fine-tuned AHMS 98.4 ± 0.9 97.9 ± 0.8 98.81 ± 1.3 98.6 ± 1.6
@@
RelCon FM From Scratch n/a 94.0 ± 1.3 92.8 ± 1.8 97.5 ± 1.1 97.0 ± 1.3
Yuan et al. (2024)’s FM From Scratch n/a 38.3 ± 12.4 23.8 ± 15.4 60.5 ± 8.6 59.6 ± 8.6
Table 3: RelCon FM compared to another Large-Scale Pre-trained Accel Model (Yuan et al., 2024). Although the RelCon FM embeds time-series into 256 dim vectors, smaller than Yuan et al. (2024)’s 1024 dim vector, it achieves stronger MLP probe performance. Additionally, fine-tuning our ResNet-34-based RelCon improves upon the from scratch performance, demonstrating the utility of our approach. Results on Yuan et al. (2024) are quoted from the original work.
Results from Comparison to a Large-Scale Pre-trained Accel Model: Table 3 shows that RelCon achieved a stronger performance across all datasets and evaluation methods. We used a ResNet-34 based architecture with an embedding dimension of 256 and 3.96M parameters, while Yuan et al. (2024) used a ResNet-18 based architecture with an embedding dimension of 1024 and 10M parameters. The smaller and deeper architecture is more effective in this case, as seen in the “From Scratch” results in Table 3. Although our learned embedding vector had a lower dimensionality, the RelCon FM has a stronger performance with the MLP probe (69.1% vs. 57.0% and 85.38% vs. 72.5% for Opportunity and PAMAP2 respectively). This suggests that the representations learned by the RelCon approach are able to efficiently capture the most important properties of the sequences. The fine-tuning evaluation shows that the initialization provided by our RelCon pre-training is able to boost performance even further with 94.0% → 98.4% and 97.5% → 98.8% improvements for Opportunity and PAMAP2 F1 metrics. This demonstrates the value of our SSL training methodology.
Results from Comparison to an Accel SSL Benchmarking Study: Table 5.2 reports results for the RelCon FM alongside results from Haresamudram et al. (2022). The results show the strong generalizability of our RelCon FM, even when there is a sensor position mismatch between training and evaluation, as seen in PAMAP2. When the target dataset’s sensor position matches the pre-training domain, as seen in HHAR, RelCon acheives exceptionally strong performance, even beating the fully-supervised methods. Although SimCLR outperforms RelCon in Motionsense, there is a considerable gap to the rest of the methods.
@@
Table 4: RelCon FM compared to the Accel SSL Benchmarking Study (Haresamudram et al., 2022). Results show the strength of the RelCon foundation model, which outperforms the fully-supervised models when the pre-training and target domains are matched to both be wrist data. In MotionSense, there is a sizable gap between SimCLR, our FM and the remaining SSL methods. Results on Haresamudram et al. (2022) are quoted from the original work.

### v5

@@ 4.1 Foundation Model Pretraining
We trained models on Inertial Movement Unit (IMU) sensor data from the Apple Heart & Movement Study (AHMS) (MacRae, 2021). AHMS is an ongoing research study designed to explore the links between physical activity and cardiovascular health, which is sponsored by Apple and conducted in partnership with the American Heart Association and Brigham and Women’s Hospital. To be eligible, participants must — among other eligibility criteria – be at least 18 years of age (at least 19 in Alabama and Nebraska; at least 21 in Puerto Rico), reside in the United States, have access to an Apple Watch, and provide informed consent electronically in the Apple Research app.
The training data included a subset of study data with 87,376 participants recorded over one day, with a 10/3/3 train/val/test split. Following prior convention (Reyes-Ortiz et al., 2015), we use 2.56 seconds of the raw 100 Hz 3-axis x,y,z accelerometry signal of the IMU sensor in the wearable device as input to our embedding model. The model was pre-trained with 8 x A100 GPUs for 24 hours. It iterated over a total of 1 billion unique samples, i.e. $\sim$30k unique participant-days. We used a 1D ResNet-34 encoder backbone (3.9M parameters) with a final global average pooling to generate a 256-dimensional embedding vector.
@@ 4.2 Downstream Evaluation
We group our downstream evaluations into two sections: Task Diversity and Benchmarking. In total, we compare against 6 different downstream datasets across 4 different types of downstream tasks. We compare our RelCon foundation model against 11 models total: 3 pre-trained from scratch in the Task Diversity Evaluation and 8 from the prior literature in the Benchmarking Evaluation.
In Task Diversity Evaluation, we evaluate how well our foundation model’s learned embedding generalizes across diverse downstream tasks, in order to understand how SSL via RelCon enables broad applicability. To this end, these downstream tasks include gait metric regression for predicting stride velocity and double support time and human activity recognition for classifying in-the-wild activities on the subsequence-level and on the workout-level. These tasks are on two separate datasets. Then, we compare our RelCon-trained foundation model against 3 different models, each pre-trained from scratch with different SOTA SSL methodology under the same, controlled conditions.
In Benchmarking Evaluation, we compare our foundation model against prior works on standardized, publicly-available human activity recognition benchmarks. Specifically, we compare against a Large-Scale Pre-trained model (Yuan et al., 2024) and against an Accel SSL Benchmarking Study (Haresamudram et al., 2022) with 7 distinct SSL methods. These datasets evaluate generalizability under data distribution shifts, including differing sensor positions and inference window lengths.
4.2.1 Task Diversity Evaluation Set-up
Baselines: Following the same exact training conditions as our RelCon foundation model (i.e. AHMS dataset, total training iterations, encoder backbone), we pre-trained 3 other state-of-the-art accelerometry SSL methods from scratch, which can seen below. Visualizations of how each method works can be found in Fig. 2. Further implementation details in Appendix A.2.
  • Accel SimCLR (Tang et al., 2020) contrasts positive pairs formed by accel-specific augmentations (i.e. 3D rotation) and demonstrates state-of-the-art performance (Haresamudram et al., 2022).
  • Augmentation Prediction (Saeed et al., 2019) predicts whether or not a specific augmentation was applied. This is a classic accelerometry approach that demonstrates competitive performance (Haresamudram et al., 2022) with good scaling ability (Yuan et al., 2024).
  • Accel REBAR (Xu et al., 2024) uses a “hard” contrastive loss with learned motif-similarity to identify positive pairs, relying on fixed positive and negative sets without capturing relative positions. For a fairer comparison, we adapt REBAR with the accel-specific enhancements from Section 3.1.
@@ 4.2.1 Task Diversity Evaluation Set-up
Gait Metric Regression: All participants completed proctored overground walking tasks with two mobile devices with IMU sensors at each side of the body in different locations (i.e. at the hip, in a front/back pocket, or in a waist bag)  (Apple, 2021). We used the Cohort A participants, which included 359 participants with an average age of 74.7, who provided informed consent. Participants conducted 3 walking tasks: 1) one lap at self-selected speed, 2) four laps at an instructed slow speed, and 3) as many laps as possible within six minutes. Each task was conducted along a straight 12-meter course, with an 8-meter pressure mat placed centrally. Various gait statistics were calculated, such as double support time and stride velocity, which our models were evaluated on.
@@
Each 2.56s subsequence was matched to the lap aggregated target (e.g. total number of steps or average walking speed). The participants were assigned into 50% train and 50% test randomly based on participant ID, where every lap for a given participant falls into the same split. The training split was used to train linear regression probes on embeddings from the self-supervised models. Metrics were selected following a prior report (Apple, 2021): mean squared error (MSE), std dev of squared error (SDSE), mean absolute error (MAE), std dev of absolute error (SDAE), and Pearson Correlation Coefficient. Mean and standard deviations were calculated by aggregating predictions across all errors of a given user and are related to bias and variance. Correlation is used to assess how each user’s average gait metric corresponds to ground truth values. Ranges per metric are obtained by retraining the probe five times and calculating the mean and standard deviation.
@@
Field Human Activity Recognition: We evaluated activity classification using self-reported activity labels gathered from in-the-wild, unproctored, field data in AHMS. We used a subset of data with $\sim$2k total users across 16 workouts (full list in Fig. 1). These workouts captured a range of diverse activities (e.g. kickboxing and rowing) that are non-trivial to separate (i.e. outdoor cycling vs. indoor cycling). For evaluation, workouts were class balanced with each including $\sim$22 hours of data, for a total of 310 workout hours. We used a 4/1/5 train/val/test split based on participant ID. We ensured this data did not overlap with the pre-training dataset, preventing any data leakage.
@@
Embeddings are generated for each 2.56s-long subsequence. We evaluate classification on both the subsequence-level as well as the workout-level. At the workout-level, we predict an activity across each of 2.56s subsequences within the workout, and select the most frequently predicted activity. Including two temporal scales which allow us to evaluate how information learned by the embedding interacts with time aggregation. We report F1, Kappa, Accuracy, and macro AUC metrics following previous work (Haresamudram et al., 2022; Yuan et al., 2024; Xu et al., 2024). Ranges per metric are obtained by retraining the probe five times and calculating the mean and standard deviation.
4.2.2 Benchmarking Evaluation Set-up
@@
Comparison to a Prior Large-Scale Pre-trained Model: Yuan et al. (2024) proposes an large-scale model with pre-trained on the UK BioBank (Bycroft et al., 2018) dataset. They are able to train over $\sim$100k unique participants for a total of $\sim$700k unique participant-days. Their model is based upon an augmentation prediction approach with a ResNet-18 backbone, and they use it to train and embed 10s subsequences of a 3-axis accelerometry signal collected at the wrist.
@@
Following their evaluation methodology, we directly compare our RelCon foundation model against their model along two different dimensions: fine-tuning the entire model and using an MLP probe. We exactly mimic their cross-validation splits on the Opportunity (Roggen et al., 2010) and PAMAP2 (Reiss, 2012) activity classification datasets and use these splits to generate the metric ranges. We also match their 10s length used for inference.
@@
Comparison to an Accel SSL Benchmarking Study: Haresamudram et al. (2022) seeks to assess the state of the accelerometry SSL field, in the context of human activity recognition. To this end, they benchmark a diverse range of self-supervised methods, including our aforementioned methods, SimCLR (Tang et al., 2020) and Augmentation Prediction (Saeed et al., 2019), as well as, SimSiam (Chen & He, 2021), BYOL (Grill et al., 2020), MAE (Haresamudram et al., 2020), CPC (Haresamudram et al., 2021), and an Autoencoder (Haresamudram et al., 2019). Each of their approaches uses a unique backbone architecture that corresponds to the original work. They pre-train their SSL methods on the 3-axis wrist-mounted accelerometry signals from Capture-24 dataset (Chan Chang et al., 2021), which is composed of 151 unique participants for a total of $\sim$4k participant-hours.
@@
Similar to our RelCon method, each of their models are pre-trained on wrist data. However, during evaluation, each activity classification dataset corresponds to accelerometry signals collected at different body positions. We compare our RelCon FM with a downstream dataset at each benchmarked position: HHAR (Blunck et al., 2015) for Wrist, Motionsense for Waist (Malekzadeh et al., 2018), and PAMAP2 (Reiss, 2012) for Leg. In this way, each method will be evaluated on potentially different sensor position as its pre-training data. We adopt the same exact cross-validation splits in the original work and use them to generate our ranges. We match their 2s length used for inference.
@@ 5.1 Task Diversity Evaluation Results
Table 1: Gait Metric Regression. Mean and Standard Deviations of Squared Error and Absolute Error were calculated by aggregating across each user and are related to bias and variance, respectively. The strong consistent performance of RelCon shows that our model is able to effectively capture the user-specific and temporally-precise information needed for gait metric regression.
Figure 3: Correlation between predicted and true DST. RelCon has the highest correlation.
Gait Metric Regression: Table 1 shows the gait metric regression performance of each SSL method.
RelCon had the strongest consistent performance across both gait metrics (1st in 3/5 evals for stride velocity and 1st in 5/5 evals for DST). While SimCLR was not specifically designed for gait analysis (Tang et al., 2020), it achieved competitive performance for stride velocity with the highest correlation. This suggests that SimCLR effectively captures overall movement patterns, but may lack the nuanced understanding required for DST, which has more complex gait dynamics compared to velocity. DST measures the brief phase when both feet are on the ground, requiring high temporal resolution to capture subtle transitions between stance and swing, influenced by neuromuscular coordination (Couto et al., 2023). In contrast, stride velocity is determined by cadence, making it simpler to model via frequency (Kirtley et al., 1985). REBAR, despite using the same learnable distance as RelCon, performed worse across most metrics due to differences in the loss functions (i.e. >.5 difference in corr for both velocity and DST). The hard contrastive loss used in REBAR relies on an absolute set of positives/negatives, limiting its capacity to capture fine-grained relational differences. In contrast, RelCon’s relative loss enabled better modeling of subtle differences of transient subject-specific gait dynamics, particularly necessary for modeling the complex neuromuscular functions that control DST, achieving the best performance across all 5 evaluations.
AHMS Activity Classification (Wrist→Wrist)
@@
Table 2: Field Human Activity Recognition across 16 Classes. RelCon achieves consistently strong performance when evaluated both at the subsequence-level and workout-level. SimCLR does well at the subsequence-level, but performs much worse at the workout-level. The opposite is true for AugPred. This shows that there exist nuances captured by our models at both levels.
Field Human Activity Recognition: Table 2 presents the activity classification performance across subseq and workout levels on real-world field data across 16 classes. Fig. 4 shows per-class performance and Fig. 5 shows detailed class predictions breakdown via a confusion matrix.
Table 2 shows that SimCLR showed strong performance at the subseq-level, with 2nd+ place across the eval metrics. However, performance dropped to 3rd- in workout-level classification. Conversely, AugPred is 2nd place in workout-level, but dropped to the worst method at the subseq-level, across all eval metrics. These differences highlight how varying temporal resolutions capture different aspects of activity dynamics, influencing model performance across time scales.
RelCon continues to achieves state-of-the-art performance with 1st place across 7/8 eval metrics. It also has the best semantic cluster separability in its embedding space, as seen in Fig. 1. As an example, Fig. 4 shows that RelCon is better at classifying stair climbing with a class AUC of .81 vs. .68 / .69 / .56. Fig. 5 shows that the other SSL methods will often confuse stair climbing with climbing or elliptical, but RelCon can capture subtle differences between the slow, deliberate hand swinging motions present in those three classes. Fig. 5 also shows that RelCon more accurately distinguishes between outdoor and indoor running (avg acc normed across the two: .69 vs. .59 / .56 / .68), likely due to its ability to detect the subtle more-uniform motion patterns characteristic of indoor running. Overall, RelCon’s strong performance across temporal resolutions and diverse activity categories underscores its robustness and generalizability in activity recognition.
Figure 4: ROC curves of Workout-Level Field Human Activity Recognition. Among other gains, RelCon is able to clearly better classify stair climbing (purple) compared to other approaches. Further class-specific prediction results can be found in the confusion matrices found in Fig. 5
@@ 5.2 Benchmarking Evaluation Results
RelCon FM MLP Probe AHMS 69.1 ± 8.3 62.1 ± 6.2 85.4 ± 3.6 84.7 ± 3.6
@@
RelCon FM Fine-tuned AHMS 98.4 ± 0.9 97.9 ± 0.8 98.8 ± 1.3 98.6 ± 1.6
@@
Table 3: RelCon FM compared to a Prior Large-Scale Pre-trained Model. Although the RelCon FM embeds time-series into 256 dim vectors, smaller than Yuan et al. (2024)’s 1024 dim vector, the frozen RelCon FM representation achieves stronger MLP probe performance.
Results from Comparison to a Prior Large-Scale Pre-trained Model: Table 3 shows that RelCon achieved a stronger performance across all datasets and evaluation methods. We used a ResNet-34 based architecture with a final embedding dimension of 256 and 3.96M parameters, while Yuan et al. (2024) used a ResNet-18 based architecture with an embedding dimension of 1024 and 10M parameters. Although our embedding vector had a lower dimensionality, RelCon FM has a stronger performance with the MLP probe (69 vs 57 / 85 vs 73 for Opportunity / PAMAP2). RelCon-learned representations are able to efficiently capture the most important properties of the sequences. Then, with fine-tuning, RelCon can achieve even stronger performance due to the strong initialization.
Results from Comparison to an Accel SSL Benchmarking Study: Table 5.2 compares the RelCon FM against a suite of SSL methods benchmarked in Haresamudram et al. (2022). Our RelCon FM shows strong generalizability, even when there is a sensor position mismatch between training and evaluation, as seen in PAMAP2. When the target dataset’s sensor position matches the pre-training domain, as seen in HHAR, RelCon acheives exceptionally strong performance, even beating the fully-supervised methods. Although SimCLR outperforms RelCon in Motionsense, there is a considerable gap between SimCLR/Relcon and the rest of the methods.
@@
Table 4: RelCon FM compared to an Accel SSL Benchmarking Study. The RelCon FM has the strongest consistent performance, even outperforming the fully-supervised model when the pre-training and target domains match, with both the pre-training dataset and downstream HHAR dataset originating from wrist sensors (Wrist→Wrist).